# Generalized Liénard systems with momentum-dependent mass: Isochronicity and bound states

### v1

@@
Generalized Liénard systems with momentum-dependent mass:
Isochronicity and bound states
Bijan Bagchi111bbagchi123@gmail.com, A. Ghose-Choudhury222aghosechoudhury@gmail.com, Aritra Ghosh333ag34@iitbbs.ac.in, Partha Guha#444partha.guha@ku.ac.ae
@@ D.H. Road, Sarisha, West Bengal 743368, India
In this paper, we explore some classical and quantum aspects of the nonlinear Liénard equation $\ddot{x}+kx\dot{x}+\omega^{2}x+(k^{2}/9)x^{3}=0$, where $x=x(t)$ is a real variable and $k,\omega\in\mathbb{R}$. We demonstrate that such an equation could be derived from an equation of the Levinson-Smith kind which is of the form $\ddot{z}+J(z)\dot{z}^{2}+F(z)\dot{z}+G(z)=0$, where $z=z(t)$ is a real variable and $\{J(z),F(z),G(z)\}$ are suitable functions to be specified. It can further be mapped to the harmonic oscillator by making use of a nonlocal transformation, establishing its isochronicity. Computations employing the Jacobi last multiplier reveal that the system exhibits a bi-Hamiltonian character, i.e., there are two distinct types of Hamiltonians describing the system. For each of these, we perform a canonical quantization in the momentum representation and explore the possibility of bound states. While one of the Hamiltonians is seen to exhibit an equispaced spectrum with an infinite tower of states, the other one exhibits branching but can be solved exactly for certain choices of the parameters.
@@ 1 Introduction
The purpose of the present investigation is to have a relook at the Liénard system (1.2) in order to highlight some intriguing facets of the problem. We will first demonstrate that generic Liénard systems (1.1) can arise from certain generalized equations of the Levinson-Smith type [13, 14]. For such Levinson-Smith equations, we will assess the feasibility of a nonlocal transformation that maps the dynamics to that of the harmonic oscillator; as a result, it will be shown that the system (1.2) can indeed be mapped to the harmonic oscillator, thus proving the isochronicity of the dynamics. Then, we will demonstrate the existence of a bi-Hamiltonian structure in the sense that there are two distinct classes of Hamiltonians which can describe the system consistently. Such Hamiltonians would turn out to be of the nonstandard kind in which there is no direct identification of kinetic and potential energies, although they admit functional forms in which the roles of the coordinate and the momentum appear interchanged with the appearance of momentum-dependent masses and potentials [9, 12]. For the derivation of such aspects, we will follow the preceding works [15, 16, 17, 18, 19, 20] which have involved the use of the Jacobi last multiplier to derive Lagrangians (and hence, Hamiltonians) for Liénard-type systems. Following this, we will proceed onto studying the quantum aspects of the system by performing a canonical quantization in the momentum representation, following [9], and employing the ordering strategy due to von Roos [21] (see [22, 23, 24, 25, 26, 27, 28, 29] for its application in systems with position-dependent masses). The time-independent Schrödinger equations for both the cases will be mapped to those of constant-mass systems but with certain effective potentials and the possibility of bound states will be investigated.
@@ 2.1 A special case
Defining $\alpha_{1}/\alpha_{2}=k$, $c_{1}=\omega^{2}$, and $c_{-1}/\alpha_{2}^{2}=k^{2}/9$, we obtain our system of interest which is (1.2).
@@ 3.2 Lagrangian and Hamiltonian functions
Let us note that not all choices for $f(x)$ and $g(x)$ would satisfy the condition [^5] For example, the van der Pol oscillator [3], i.e., $\ddot{x}+\sigma(1-x^{2})\dot{x}+\Omega_{0}^{2}x=0$ for $\sigma,\Omega_{0}>0$ does not satisfy (3.7). (3.7), often dubbed the Chiellini integrability condition in the context of Abel equations of the first kind [33]. Thus, Liénard systems for which the functions $f(x)$ and $g(x)$ satisfy the Chiellini condition, (1.1) may be rewritten as (3.6). If that is the case, then combining (3.5) with (3.6) gives us the following expression for the Jacobi last multiplier:
@@ 3.3 Hamiltonians corresponding to the system of our interest
It is at once apparent that the above Lagrangian (3.10) is of the nonstandard form in that it involves a complicated expression of the derivative rather than being controlled by the usual kinetic-energy term which is quadratic in velocity. Of late, such non-conventional forms of Lagrangians have been a subject of active pursuance in the literature (see [12] for a recent review). These Lagrangians often lead to certain exotic Hamiltonians, as typified by the so-called branched Hamiltonians [37, 38], whose connections to problems of nonlinear dynamics, as governed by autonomous differential equations, have been well researched [39, 40]. We will distinguish the two solutions of $\ell$ furnished in (3.15) as ‘class-I’ and ‘class-II’ solutions, respectively, for $\ell=-1/3,-2/3$.
@@ 4.2 Class-II
where $\epsilon$ is as defined in the line below equation (4.9). Once again, we will perform the transformation
@@ 5.1 Class-II
For this case, the effective potential has been plotted in Fig. (3) and clearly shows the isotonic behavior [56, 57, 58]. The wavefunctions can be computed straightforwardly and involve the confluent hypergeometric functions (see Appendix (B)) which are in-turn related to Laguerre polynomials [58, 59]. The spectrum consists of an infinite tower of states with equispaced energy levels. Defining $g=\epsilon-\frac{1}{4}+\frac{96}{k}$, the effective potential $V_{\rm II}|_{\rm eff}(\zeta)$ reads
@@
For all values of $\epsilon\geq 0$ [24, 25, 26, 27, 28, 29], the spectrum (level spacing) remains the same although the ground-state energy is dependent on $\epsilon$ and $k$. Then, up to normalization factors, the eigenfunctions of (4.19) turn out to be (see Appendix (B))
@@ 5.2 Class-I
$V^{\pm}_{\rm I}|_{\rm eff}(\xi)=\frac{\omega^{2}}{64}(\xi\mp\xi_{0})^{2}+\frac{\Delta}{\xi^{2}}-\frac{k}{24\omega^{2}},\quad\quad\xi_{0}=\sqrt{\frac{k}{6}}\frac{4}{\omega^{2}},\quad\Delta=\epsilon-\frac{1}{4},$ (5.4)
@@
where $\xi>0$. While these naively appear to be identical to the harmonic potential (with center at ±$\pm\xi_{0}$) for a particle of mass $m=1/2$ and frequency $\omega/4$, there is a peculiarity because $\xi>0$ and this must be accounted for. Explicitly, the necessary condition for the existence of bound states is as follows –
@@
is a bound state of the potential (5.4) $V^{\pm}_{\rm I}|_{\rm eff}(\xi)$ for $\Delta=0$, if and only if
@@ 6 Conclusions
While the classical Hamilton’s equations are equivalent for both the Hamiltonians (when they are coupled to give rise to the second-order equation for the coordinate variable), we demonstrated that their quantum-mechanical features are largely different. One of the Hamiltonians which we denoted earlier to be of class-II admits an infinite tower of equispaced energy levels, giving us a nontrivial example of the correspondence between isochronicity and equispaced quantum spectra [8]. Further, we showed that the class-I scenario admits branching, leading to a pair of branched Hamiltonians, both of which exhibit peculiarities with the boundary conditions which in-turn obstructs the presence of an infinite tower of equispaced levels.
@@
Acknowledgements
B.B. and A.G. would like to thank Miloslav Znojil for discussions and for collaboration on [12]. A.G. expresses his gratitude to the Ministry of Education, Government of India for financial support in the form of a Prime Minister’s Research Fellowship (ID: 1200454) and also thanks the Czech Technical University in Prague for hospitality where many of the ideas presented in this work were conceived. A.G. also thanks Akash Sinha for various related discussions. A.G.C. and P.G. express their gratitude to Pepin Cariñena for stimulating discussions. We thank Jasleen Kaur for help in preparing the figures and for carefully reading the manuscript.

### v2

# Generalized Liénard systems and isochronous connections

@@
Generalized Liénard systems and isochronous connections
Bijan Bagchi111bbagchi123@gmail.com, A. Ghose-Choudhury222aghosechoudhury@gmail.com, Aritra Ghosh333ag34@iitbbs.ac.in; aritraghosh500@gmail.com, Partha Guha#444partha.guha@ku.ac.ae
@@ D.H. Road, Sarisha, West Bengal 743368, India
In this paper, we explore some classical and quantum aspects of the nonlinear Liénard equation $\ddot{x}+kx\dot{x}+\omega^{2}x+(k^{2}/9)x^{3}=0$, where $x=x(t)$ is a real variable and $k,\omega\in\mathbb{R}$. We demonstrate that such an equation could be derived from an equation of the Levinson-Smith kind which is of the form $\ddot{z}+J(z)\dot{z}^{2}+F(z)\dot{z}+G(z)=0$, where $z=z(t)$ is a real variable and $\{J(z),F(z),G(z)\}$ are suitable functions to be specified. It can further be mapped to the harmonic oscillator by making use of a nonlocal transformation, establishing its isochronicity. Computations employing the Jacobi last multiplier reveal that the system exhibits a bi-Hamiltonian character, i.e., there are two distinct types of Hamiltonians describing the system. For each of these, we perform a canonical quantization in the momentum representation and explore the possibility of bound states. While one of the Hamiltonians is seen to exhibit an equispaced spectrum with an infinite tower of states, the other one exhibits branching but can be solved exactly in closed form for certain choices of the parameters.
@@ 1 Introduction
The purpose of the present investigation is to have a relook at the Liénard system (1.2) in order to highlight some intriguing facets of the problem. We will first demonstrate that generic Liénard systems (1.1) can arise from certain generalized equations of the Levinson-Smith type [13, 14]. For such Levinson-Smith equations, we will assess the feasibility of a nonlocal transformation that maps the dynamics to that of the harmonic oscillator; as a result, it will be shown that the system (1.2) can indeed be mapped to the harmonic oscillator, thus proving the isochronicity of the dynamics. Then, we will demonstrate the existence of a bi-Hamiltonian structure in the sense that there are two distinct classes of Hamiltonians which can describe the system consistently. Such Hamiltonians would turn out to be of the nonstandard kind in which there is no direct identification of kinetic and potential energies, although they admit functional forms in which the roles of the coordinate and the momentum appear interchanged with the appearance of momentum-dependent masses and potentials [9, 12]. For the derivation of such aspects, we will follow the preceding works [15, 16, 17, 18, 19, 20] which have involved the use of the Jacobi last multiplier to derive Lagrangians (and hence, Hamiltonians) for Liénard-type systems. Following this, we will proceed onto studying the quantum aspects of the system by performing a canonical quantization in the momentum representation, following [9] and employing the ordering strategy due to von Roos [21] (see [22, 23, 24, 25, 26, 27, 28, 29] for its application in systems with position-dependent masses). The time-independent Schrödinger equations for both the cases will be mapped to those of constant-mass systems but with certain effective potentials and the possibility of bound states will be investigated.
@@ 2.1 A special case
Defining $\alpha_{1}/\alpha_{2}=k$, $c_{1}=\omega^{2}$, and $c_{-1}/\alpha_{2}^{2}=k^{2}/9$, we can obtain our system of interest which is (1.2).
@@ 3.2 Lagrangian and Hamiltonian functions
Let us note that not all choices for $f(x)$ and $g(x)$ would satisfy the condition [^5] For example, the van der Pol oscillator [3], i.e., $\ddot{x}+\sigma(1-x^{2})\dot{x}+\Omega_{0}^{2}x=0$ for $\sigma,\Omega_{0}>0$ does not satisfy (3.7). (3.7), often dubbed the Chiellini integrability condition in the context of Abel equations of the first kind [33]. Thus, for Liénard systems for which the functions $f(x)$ and $g(x)$ satisfy the Chiellini condition, (1.1) may be rewritten as (3.6). If that is the case, then combining (3.5) with (3.6) gives us the following expression for the Jacobi last multiplier:
@@ 3.3 Hamiltonians corresponding to the system of our interest
It is at once apparent that the above Lagrangian (3.10) is of the nonstandard form in that it involves a complicated expression of the derivative rather than being controlled by the usual kinetic-energy term which is quadratic in velocity. Of late, such non-conventional forms of Lagrangians have been a subject of active pursuance in the literature (see [12] for a recent review). These Lagrangians often lead to certain exotic Hamiltonians, as typified by the so-called branched Hamiltonians [37, 38], whose connections to problems of nonlinear dynamics, as governed by autonomous differential equations, have been well researched [39]. We will distinguish the two solutions of $\ell$ furnished in (3.15) as ‘class-I’ and ‘class-II’ solutions, respectively, for $\ell=-1/3,-2/3$.
@@ 4.2 Class-II
where $\epsilon$ is defined in the line below (4.9). Once again, we will perform the transformation
@@ 5.1 Class-II
For this case, the effective potential has been plotted in Fig. (3) and clearly shows the isotonic behavior [55, 56, 57]. The wavefunctions can be computed straightforwardly and involve the confluent hypergeometric functions (see Appendix (B)) which are in-turn related to the associated Laguerre polynomials [57, 58]. The spectrum consists of an infinite tower of states with equispaced energy levels. Defining $g=\epsilon-\frac{1}{4}+\frac{96}{k}$, the effective potential $V_{\rm II}|_{\rm eff}(\zeta)$ reads
@@
For all values of $\epsilon\geq 0$ [24, 25, 26, 27, 28, 29], the spectrum (level spacing) remains the same although the ground-state energy is dependent on $\epsilon$ and $k$. Up to normalization factors, the eigenfunctions of (4.19) turn out to be (see Appendix (B))
@@ 5.2 Class-I
$V^{\pm}_{\rm I}|_{\rm eff}(\xi)=\frac{\omega^{2}}{64}(\xi\mp\xi_{0})^{2}+\frac{\Delta}{\xi^{2}}-\frac{k}{24\omega^{2}},\quad\quad\xi_{0}=\sqrt{\frac{k}{6}}\frac{4}{\omega^{2}},\quad\quad\Delta=\epsilon-\frac{1}{4},$ (5.4)
@@
where $\xi>0$. While these naively appear to be identical to the harmonic potential (with center at ±$\pm\xi_{0}$) for a particle of mass $m=1/2$ and frequency $\omega/4$, there is a peculiarity because $\xi>0$ and this must be accounted for. Explicitly, a sufficient condition for the existence of bound states is as follows –
@@
is a bound state of the potential (5.4) $V^{\pm}_{\rm I}|_{\rm eff}(\xi)$ for $\Delta=0$, if
@@ 6 Conclusions
While the classical Hamilton’s equations are equivalent for both the Hamiltonians (when they are coupled to give rise to the second-order equation for the coordinate variable), we demonstrated that their quantum-mechanical features are largely different. One of the Hamiltonians which we denoted earlier to be of class-II admits an infinite tower of equispaced energy levels, giving us a nontrivial example of the correspondence between isochronicity and equispaced quantum spectra [8]. Further, we showed that the class-I scenario admits branching, leading to a pair of branched Hamiltonians, both of which exhibit peculiarities with the boundary conditions.